Speedy Gonzales and the Road Runner are racing each other in the 100 m long dash. Speedy travels at a constant velocity of 6.80 m/s for the entire race and the Road Runner accelerates 2 mfrom rest at 1.60 m/s . Who wins the race? 


1
Expert's answer
2021-01-14T10:40:50-0500
"t_1=\frac{s}{v}=\frac{100}{6.8}=14.7\ s"

"t_2=\sqrt{2\frac{s}{a}}=\sqrt{2\frac{100}{1.6}}=11.2\ s<t_1"

Road Runner wins the race.